# Spatio-Temporal Graphical Counterfactuals: An Overview

### v1

@@ Abstract
Counterfactual thinking is a critical yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve their performances for new scenarios. Many research works, including Potential Outcome Model and Structural Causal Model, have been proposed to realize it. However, their modelings, theoretical foundations and application approaches are usually different. Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units. Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal frameworks to infer the spatio-temporal counterfactuals.
Index Terms:
@@
Counterfactual inference, spatio-temporal graphical model, Potential Outcome Model, Structural Causal Model, causality.
I Introduction
@@
How to enable an intelligent machine to think and answer a counterfactual question? E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns? Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Or, would someone still purchase the product even if they had never been shown the advertisement? These questions generally have the counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”. Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past. Secondly, counterfactual outcomes are unobservable. The outcomes of some actions are observed if the actions have been done to the real-world system, thus these outcomes are factual. But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done.
Figure 1: Infer counterfactuals with less and less knowledge.
To enable an intelligent machine to imitate counterfactual thinking like human, there are mainly three approaches, as shown in Fig. 1. The first is system identification, which aims to identify the system dynamics, and then solve the evolution trajectories with different initial values [1]. This process may be data-driven and adaptive, but the physical mechanisms and models are necessary to be known as knowledge in prior [2, 3, 4, 5]. E.g., in [6, 7], functional basis or network structure are necessary to be known at least one to identify the network dynamics sparsely. However, the knowledge of system mechanism is usually hard to obtain in prior.
The second approach is generative modeling, which aims to model and learn the system uncertainty, and then generate scenarios with different random inputs [8, 9]. Thus, compared with system identification, generative approach is not necessary to know the determined part, that is the knowledge of system mechanism, but on contrary, it focuses on the uncertain modeling, especially with the deep learning [10, 11, 12, 13, 14, 15]. However, the generated scenarios would not be out of the distribution of the observational data [9], while the counterfactual scenarios are usually not in the observational factuals. Thus, the mechanism knowledge is still needed to correct the bias between the distributions of factuals and counterfactuals.
Thus, the third approach is causal inference, which aims to discover causality from the observational data, and then infer factuals and counterfactuals directly with as little knowledge as possible [16, 17]. Causal counterfactual inference enables artificial intelligence to learn the knowledge from data, and ultimately improve their performances for new scenarios. To realize it, many research works have been proposed to infer counterfactuals, including the temporal cases [18, 19, 20, 21, 22, 23]. However, there are still two challenging but critical issues to address: (i) Most current inference approaches are built on the frameworks of Potential Outcome Model (POM) [17] or Structural Causal Model (SCM) [16, 24, 25, 26]. The goals of them are the same, but their foundations are different. This induces difficulty in discussing them in a unified framework. (ii) The two frameworks both cannot infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between system units.
Thus, the main contributions of this work are as follows:
@@
    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section II to Section IV.
@@
    A Forward Counterfactual Inference algorithm is designed here to recursively and autonomously infer counterfactuals through causal graphical languages. This part is presented as Algorithm 1 in Section IV-C.
@@
    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks. This part is organized in Section V.
II Potential Outcome Model
The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in TABLE I. Here, $Y$ is the potential outcomes of the target variable. $T$ is treatment variable. $T=1$ represents being treated and $T=0$ represents being controlled, with a group of randomized control trials on $N$ units. $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials. POM models the general process from causal estimands ($T$, $Y$ and $X$) to “Science” [27]. In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
TABLE I: Data for POM example.
@@ 5 $x_{5}$ $y_{5}(1)$??? $1$
$N$ $x_{N}$??? $y_{N}(0)$ 00
II-A Inferring Counterfactuals through POM
@@
Another common approach is data-driven imputation, that views the counterfactuals as missing values, and interpolates them by fitting the observational data. To realize it, linear or nonlinear regressions [36, 37, 38, 39] are commonly used to interpolate the missing values from the trained regressive models on the observational data. Different from this, tensor decomposition approaches [40, 41, 42] recover the missing values by decomposing and reconstructing the sparse data tensor. Moreover, deep generative models [10, 11, 12, 13, 14, 15] are also used to model the uncertainty in data, and generate the missing values by randomly sampling. Note that, the generated missing values are not unique, but a group of values conforming to a random distribution. This is different compared to the regressive approaches and tensor decomposition approaches.
II-B Foundational Assumptions of POM
Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is the stable unit treatment value assumption (SUTVA), defined as
Assumption 1 (SUTVA [43]).
In TABLE I,
@@ II-B Foundational Assumptions of POM
The second assumption is consistency, defined as
Assumption 2 (Consistency [44, 45]).
In TABLE I, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units.
Another form to describe the consistency assumption through expectation $\mathbb{E}[\cdot]$ is
@@
where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in TABLE I. This is the same to the case of conditional expectation with $X$. Thus, the consistency assumption guarantees the missing counterfactual values in Section II-A can be interpolated by the other observational values actually.
The third assumption is positivity, defined as
Assumption 3 (Positivity [32]).
In TABLE I, $10<P(T=1|X=x)<$ 1.
This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$. Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
The forth assumption is ignorability, defined as
Assumption 4 (Ignorability [44, 45]).
In TABLE I, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit.
Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$. Thus, the ignorability assumption is also called as unconfoundedness. This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [46].
III Structural Causal Model
An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$. Note that, endogenous variables are observable from data, but exogenous variables are not. Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ are also set to describe the functional relationship between these variables.
@@ III Structural Causal Model
Figure 3: A diagram of Pearl’s causal ladder [24].
III-A Pearl’s Causal Ladder [16, 24, 26]
@@
    Action: Perform an intervention on SCM, e.g., $do(X_{2}=2)$ on SCM in Fig. 2, and then, $X_{2}$ would be assigned with value $2$, and the arrows to $X_{2}$ would be modified, as shown in Fig. 3;
@@
The intervention, also named as $do$-operator [16], can be generalized into the form of probability, if modularity assumption is introduced as follow:
Assumption 5 (Modularity [16]).
If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
@@
III-B Foundational Assumptions of Bayesian Network
@@
Then, two assumptions, causal Markov (or Markov property) and faithfulness, are introduced as follows:
Assumption 6 (Causal Markov [47, 48, 49, 50, 51]).
Probability distribution $P$ is Markovian to a Bayesian network $\mathcal{G}$ if
@@
where $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ are three disjoint subsets.
Assumption 7 (Faithfulness [47, 48, 49, 50, 51]).
Probability distribution $P$ is faithful to a Bayesian network $\mathcal{G}$ if
@@
for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$.
It is intuitive to understand the two assumptions, that is, the Bayesian network $\mathcal{G}$ has one-to-one correspondence to the independence of probability distribution $P$ in observational data. Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follow:
@@
Moreover, a Bayesian network is also assumed to be causally sufficient, that is
Assumption 8 (Causal Sufficiency [47, 48, 49, 50, 51]).
Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$.
@@
This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network. Actually, the ignorability assumption in POM (see Assumption 4) also implies the causal sufficiency, and this assumption is weaker than ignorability. But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually. Thus, some research works focus on this issue. For example, FCI (or Fast Causal Inference) algorithm and its variants [66, 67, 68, 69, 70, 71] are proposed to discover causality in the presence of hidden variables. And some other related works [72, 73, 74] investigate the causal discovery with soft intervention, that intervenes the SCM but does not change the network structure. There are also some related works [75, 76, 77, 78] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.
IV Differences between POM and SCM
Through the above comparisons between POM and SCM, one can find many differences intuitively. For example, they have different statements for counterfactuals, even if some of them can be equally transformed for each other. Moreover, SCM has a causal graph to represent the causations, and the interventions are also introduced to modify the graph to obtain the counterfactuals, but POM does not use it. However, although they have made many distinctions to distinguish each other, as stated in Rubin’s and Pearl’s books [17, 26], there are still some fundamental differences and relationship that need to be discussed.
IV-A Ignorability and Back-door Criterion

### v2

# Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview

@@ Abstract
Counterfactual thinking is a critical yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve their performances for new scenarios. Many research works, including Potential Outcome Model and Structural Causal Model, have been proposed to realize it. However, their modelings, theoretical foundations and application approaches are usually different. Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units. Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal framework to infer the spatio-temporal counterfactuals.
keywords:
@@
Counterfactual inference, spatio-temporal graphical model, Potential Outcome Model, Structural Causal Model, causality
1 Introduction
@@
How to enable an intelligent machine to think and answer a spatio-temporal counterfactual question? E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns? Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Or, would someone still purchase the product even if they had never been shown the advertisement? These questions generally have the spatio-temporal counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”. Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past. Secondly, counterfactual outcomes are unobservable. The outcomes of some actions are observed if the actions have been done to the real-world system, thus these outcomes are factual. But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done. Thirdly, the real-world system can be observed as multivariate time series, and there are widespread multivariate interactions within the time series [1, 2].
To enable an intelligent machine to imitate counterfactual thinking like human, thus, Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] are proposed as two frameworks with different foundations. POM uses stable unit treatment value assumption (SUTVA), consistency assumption, positivity assumption and ignorability assumption as foundations [8, 9, 10, 11]. If these assumptions are satisfied, one can use matching or imputation approaches to infer counterfactuals without confounders. But in contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7]. With the graphical language d-separation, then causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
However, the different foundations between POM and SCM lead to differences in counterfactual inference. Such differences include: (i) POM assumes the positivity, but SCM does not. That means the counterfactuals in POM are factual actually and can be found in observations certainly. But in contrast, the counterfactuals in SCM can be out of observations and unfalsifiable, that is, one cannot validate the correctness of the inferred counterfactuals. (ii) POM assumes the consistency, but it is rejected for SCM [7]. Actually, SCM also needs the consistency assumption implicitly to guarantee that the counterfactual outcome is one of the unobserved potential outcomes. (iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, that makes it difficult to satisfy the ignorability. But in contrast, SCM can deconfound through back-door criterion [17, 18, 4, 5] based on the casual sufficiency assumption, and it is equivalent to the ignorability actually. (iv) It is difficult for POM to analyze indirect causation (mediation), thus POM typically relies on Baron-Kenny model [19, 20] for assistance. By contrast, it is easier for SCM by using front-door criterion [17, 18, 4, 5], that decomposes a causal pathway into two parts, i.e., from source variable to mediator and from mediator to target variable, and then conduct back-door adjustment twice if no other mediator.
Moreover, whether it is POM or SCM currently, they are both unable to model the spatio-temporal interactions between system units. E.g., the SUTVA requires no interaction between system units in POM. Moreover, SCM based on the traditional Bayesian network has no modeling for multivariate time series, thus the nonstationarity in time series is often neglected. Thus, this motivates us to find a common point to unify the two frameworks, and strengthen the inference for the spatio-temporal counterfactuals. Fortunately, our recent work [21, 22] has proposed a Spatio-Temporal Bayesian Networks (STBNs) model based on temporal assumption, that assumes the cause precedes or parallels the effect in time. And it is found that, the famous complex networks [23, 24, 25], a type of network dynamics to model spatio-temporal interactions, are a special case of STBNs.
Thus, we further conduct this work to build a spatio-temporal graphical counterfactual framework based on STBNs. The framework requires two components: (i) an STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion. Then we discuss how to get the two points, following the storyline as shown in Fig. 1. Thus, the main contributions of this work are as follows:
@@
    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
@@
    A Forward Counterfactual Inference algorithm is designed here to recursively and autonomously infer counterfactuals through causal graphical languages. This part is presented as Algorithm 1 in Section 4.3.
@@
    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks. This part is organized in Section 5.
Figure 1: Towards spatio-temporal graphical counterfactuals.
2 Potential Outcome Model
The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in Table 1. Here, $Y$ is the potential outcomes of the target variable. $T$ is treatment variable. $T=1$ represents being treated and $T=0$ represents being controlled, with a group of randomized control trials on $N$ units. $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials. POM models the general process from causal estimands ($T$, $Y$ and $X$) to “Science” [26]. In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
Table 1: Data for POM example.
@@ 5    $x_{5}$    $y_{5}(1)$    ??    11
   NN    $x_{N}$    ??    $y_{N}(0)$    0
2.1 Inferring Counterfactuals through POM
@@
Another common approach is data-driven imputation, that views the counterfactuals as missing values, and interpolates them by fitting the observational data. To realize it, linear or nonlinear regressions [34, 35, 36, 37] are commonly used to interpolate the missing values from the trained regressive models on the observational data. Different from this, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing the sparse data tensor. Moreover, deep generative models [41, 42, 43, 44, 45, 46, 47] are also used to model the uncertainty in data, and generate the missing values by randomly sampling. Note that, the generated missing values are not unique, but a group of values conforming to a random distribution. This is different compared to the regressive approaches and tensor decomposition approaches.
2.2 Foundational Assumptions of POM
Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
@@ 2.2 Foundational Assumptions of POM
The second assumption is consistency, defined as {assumption}[Consistency [8, 9]] In Table 1, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units. Another form to describe the consistency assumption through expectation $\mathbb{E}[\cdot]$ is
@@
where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in Table 1. This is the same to the case of conditional expectation with $X$. Thus, the consistency assumption guarantees the missing counterfactual values in Section 2.1 can be interpolated by the other observational values actually.
The third assumption is positivity, defined as {assumption}[Positivity [11]] In Table 1, $0<P(T=1|X=x)<1$. This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$. Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
The forth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit. Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$. Thus, the ignorability assumption is also called as unconfoundedness. This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [48].
3 Structural Causal Model
Different from POM, SCM provides a type of graphical language to discuss causality intuitively, which avoids complex human-guided technical operations. An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$. Note that, endogenous variables are observable from data, but exogenous variables are not. Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ are also set to describe the functional relationship between these variables.
@@ 3 Structural Causal Model
3.1 Pearl’s Causal Ladder [4, 5, 7]
@@
    Action: Perform an intervention on SCM, e.g., $do(X_{2}=2)$ on SCM in Fig. 2, and then, $X_{2}$ would be assigned with value 22, and the arrows to $X_{2}$ would be modified, as shown in Fig. 3;
@@
The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if modularity assumption is introduced as follow: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
@@
3.2 Foundational Assumptions of Bayesian Network
@@
Then, two assumptions, causal Markov (or Markov property) and faithfulness, are introduced as follows: {assumption}[Causal Markov [12, 13, 14, 15, 16]] Probability distribution $P$ is Markovian to a Bayesian network $\mathcal{G}$ if
@@
where $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ are three disjoint subsets. {assumption}[Faithfulness [12, 13, 14, 15, 16]] Probability distribution $P$ is faithful to a Bayesian network $\mathcal{G}$ if
@@
for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$. It is intuitive to understand the two assumptions, that is, the Bayesian network $\mathcal{G}$ has one-to-one correspondence to the independence of probability distribution $P$ in observational data. Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follow:
@@
Moreover, a Bayesian network is also assumed to be causally sufficient, that is {assumption}[Causal Sufficiency [12, 13, 14, 15, 16]] Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$. This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network. Actually, the ignorability assumption in POM (see Assumption 2.2) also implies the causal sufficiency, and this assumption is weaker than ignorability. But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually. Thus, some research works focus on this issue. For example, FCI (or Fast Causal Inference) algorithm and its variants [61, 62, 63, 64, 65, 66] are proposed to discover causality in the presence of hidden variables. And some other related works [67, 68, 69] investigate the causal discovery with soft intervention, that intervenes the SCM but does not change the network structure. There are also some related works [70, 71, 72, 73] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.
4 Differences between POM and SCM
In this section, the fundamental differences between POM and SCM are discussed, including (i) counterfactual falsifiability: POM assumes the positivity, thus the counterfactuals in POM can be falsified on observations, but SCM does not. The counterfactuals in SCM can be outside the observations, thus it is unfalsifiable. (ii) using back-door criterion to satisfy ignorability: POM requires ignorability to guarantee the unconfoundedness and correctness of counterfactual inference, but it is a complex human-guided operation. By contrast, SCM realizes unconfoundedness by the causal sufficiency and back-door adjustment, that are equivalent to the ignorability assumption. SCM facilitates counterfactual inference through graphical language. (iii) direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on Baron-Kenny model. But in contrast, SCM has advantages to adjust the front-door path (that has indirect causation) through equivalently applying the back-door criterion twice. We further use this advantage to design the forward counterfactual inference algorithm to infer the spatio-temporal counterfactuals autonomously. The following are the details.
4.1 Counterfactual Falsifiability
As shown in Fig. 3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$. That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 2.2). And then, due to the consistency assumption (see Assumption 2.2), we can infer the counterfactuals through matching the two samples exactly. But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1. Thus, how to falsify the counterfactuals?
It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in Table 1), and we need at least two samples for comparison. In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times. Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [7].
Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable. For example, as shown in Fig. 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$. Then, we query “if had done $T=4$, what would Y=?Y=?”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation. Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations. This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
4.2 Ignorability and Back-door Criterion